OpenStudy (anonymous):

7c4/7p4 evaluate the expression

OpenStudy (anonymous):

if you mean 7c^4/7p^4, then it'd just be c^4/p^4. Otherwise, I have no idea lol.

OpenStudy (anonymous):

by the combination

OpenStudy (m):

7c4 = 7!/(7-4)!4!

OpenStudy (m):

7p4 = 7!/(7-4)!

OpenStudy (anonymous):

7!(3!)4!

OpenStudy (m):

yeah

OpenStudy (m):

for the bottom

OpenStudy (m):

calculate those two seperately

OpenStudy (anonymous):

so u do 7*6*5 then 4*3*2

OpenStudy (m):

7*6*5*4/4*3*2*1

OpenStudy (m):

for 7c4 = 7*6*5*4

OpenStudy (m):

i mean 7p4

OpenStudy (anonymous):

840/24 =35

OpenStudy (m):

1260/840 = 1.465

OpenStudy (m):

whoops it's 35/840

OpenStudy (anonymous):

so id leave it as a decimal like that?

OpenStudy (m):

or fraction

OpenStudy (anonymous):

but if u do 35/840 it goes into 24

OpenStudy (m):

it's .04166

OpenStudy (anonymous):

.04166 for the whole answer?

OpenStudy (m):

yeah

OpenStudy (anonymous):

because it says leave it as a fraction or an integer

OpenStudy (m):

35/840 or .04166

OpenStudy (m):

it's the correct answer :)

OpenStudy (m):

fraction or an integer is a contradiction

OpenStudy (m):

your teacher meant fraction or decimal

OpenStudy (anonymous):

so leave it as a fraction or put the .04166

OpenStudy (m):

either is fine but fraction is better because it's more accurate

OpenStudy (m):

use fraction

OpenStudy (anonymous):

Ok I just have one last similar one